Calculate the number of molecules that are in 69.2 grams of dihydrogen monoxide
1
Expert's answer
2020-05-12T11:34:52-0400

N(molecules H2O)=n*N Avogadro

n(H2O)=m(H2O)/Mr(H2O)=69,2/18=3,84

N(molecules H2O)=3,84*(6*10^23)=2,304*10^24 molecules